# Quantum phase estimation based filtering: performance analysis and application to low-energy spectral calculation

### v1

@@ II.4 The Gibbs phenomenon and the $\sigma$-factor
As seen in the previous section, the QPE-based filter $\mathcal{R}_{\mu}$ generally exhibits oscillation known as the Gibbs phenomenon [43]. To analyze this effect, let us first rewrite $\mathcal{R}_{\mu}$ as
@@
In the theory of the Gibbs phenomenon [43], this $\sigma$-factor is called a “filter” (not to be confused with filtering of states discussed in earlier sections) and serves as a smoothing factor that improves the convergence of the truncated Fourier series. Several filters have been proposed in the classical Fourier spectral analysis. Some examples are the Lanczos filter and the raised cosine filter [43, 44].
@@ II.5 Comparison with QETU
In this section, we compare the QPE-based and the QETU-based filters. Figure 7 shows the quantum circuit for QETU [12] with one ancilla qubit. It consists of $d+1$ rotation gates with symmetric $d+1$ phase angles $(\varphi_{0},\varphi_{1},\dots,\varphi_{1},\varphi_{0})$ and $d$ blocks of $U^{\dagger}_{HT/2}=e^{-iHT/2}$ and $U_{HT/2}$ (note that in our case $HT$ is defined in the interval $[0,2\pi]$). The output of this circuit after measuring the ancilla qubit is $\frac{f_{d}(w)|\psi\rangle}{||f_{d}(w)|\psi\rangle||}$, where $f_{d}(w=\cos\frac{HT}{4})$ is an even polynomial of degree $d$. It has been shown in Ref. [12] that the QETU-based filter can be constructed with $d=O(\delta^{-1}\log\epsilon^{-1})$ operations of $U^{\dagger}_{HT/2}$ and $U_{HT/2}$. This is the same order as the Kaiser window-based QPE filter, although in the case of the Kaiser window the extra cost of the window function preparation needs to be taken into account. For example, in the approach of Ref. [42], a system-independent cost of $O(n\sqrt[4]{\alpha+1}(\alpha+\log\epsilon^{-1}))=O(n(\log\epsilon^{-1})^{5/4})$ has to be added.
@@
where $T_{2l}(w)$ and $c_{2l}$ are the $2l$-th Chebyshev polynomials and expansion coefficients, respectively. We numerically determine the smallest $d$ in which $F_{d}(E)=|f_{d}(\textrm{cos}\frac{ET}{4})|^{2}$ satisfies Eqs. (42-44) for all $E\in[0,\frac{2\pi}{T}]$. The expansion coefficients $c_{2l}$ are calculated with the convex-optimization-based approach proposed in Ref. [12] and using the qsppack library [45, 46, 47]. We use $M$ grid points $w_{m=0,1,\dots,M-1}$ in $[0,1]$ and solve the following optimization problem
@@
After the optimal coefficients $c_{2l}$ are obtained with $M$ sampled grid points, we check whether the function $F_{d}(E)$ satisfies Eqs. (42-44) for all $E\in[0,\frac{2\pi}{T}]$ by combining a grid search with a fine grid of $100,000$ points and an optimization routine using the scipy.optimization module [48]. As an example of the calculated results, $f_{d}(w)$ and $F_{d}(E)$ obtained with degree $144$ and $\epsilon=0.01$ are shown in Figs. 8(a) and 8(b).
@@
Figure 8: Polynomial approximations of (a) $f(w)$ and (b) $F(E)=|f(\cos\frac{ET}{4})|^{2}$ with degree $144$ and $\epsilon=0.01$. The parameters used are $E_{\textrm{targ}}=1.4430,2\delta=0.2556$, and $T=1$. The two vertical dotted lines indicate the positions of $w_{\pm}$ in (a) and $E_{\textrm{targ}}$ and $E_{\textrm{targ}}+2\delta$ in (b). The inset in (b) shows a magnified view.
@@
Figure 9: Minimum number of queries to $U_{HT}$ in QETU obtained with $M=800$, $3200$, and $12800$ for three sets of ($E_{\textrm{targ}},\delta$). The corresponding values in the QPE-based filters are also shown for comparison. The straight lines show the results of the linear regression. (a) $E_{\textrm{targ}}=1.0589$ and $1.02322\delta=$ 1.0232. (b) $E_{\textrm{targ}}=1.4430$ and $0.25562\delta=$ 0.2556. (c) $E_{\textrm{targ}}=1.5069$ and $0.12792\delta=$ 0.1279. In (c) the results for $\epsilon=10^{-4}$ and $\epsilon=10^{-5}$ with $M=800$ are not calculated as there are less sampled points than the required degree $d$.
@@
Figure 9 shows the calculated number of queries to $U_{HT}$, which is half of the minimum degree $d$, for three different sets of $E_{\textrm{targ}},\delta$. Here $E_{\textrm{targ}}$ and $\delta$ are obtained numerically from the results of the Kaiser window-based filter (Fig. 5 (b)) with $\alpha=3$ and $n=6,8,9$, using $y_{c}=2^{n-2}-1$ and $\epsilon=10^{-7}$ (see also Fig. 4). The number of sampled points $M$ is a hyperparameter, and we compare three cases, $M=800,3200$ and $12800$. In the figures, $\epsilon$ is varied from $10^{-2}$ to $10^{-5}$. For $\epsilon<10^{-5}$ we have difficulty obtaining reliable results. The errors in the QPE-based filters with the three window functions are also shown for comparison.
@@
In QETU the number of queries increases linearly with $\log(\epsilon^{-1})$ as expected, and decreases as the number of sampled points $M$ increases. The latter trend can be understood by noting that with a small $M$ the optimized coefficients tend to overfit to the sampled data points. Interestingly, for the three cases studied, the QETU results extrapolated to $\epsilon=10^{-7}$ are very close to the results of the Kaiser window-based filter (crosses). This can potentially be understood by noting that the Kaiser window is an approximation of the prolate spheroidal window, which is designed to maximize the probability of the mainlobe of the probability distribution $P_{\mu}(y)$ [26]. This construction is approximately equivalent to minimizing the error of the QPE-based filter $\Delta\mathcal{R}_{\mu}$ defined in Eq. (14), potentially explaining the similarities between the QETU-based filter obtained from the optimization (Eq. (52)) and the Kaiser window-based filter. Although the extra cost of embedding the window function $a_{j}$ and $O(\log\bigl{[}\delta^{-1}\log\epsilon^{-1}\bigr{]})$ ancilla qubits may prevent the Kaiser window-based filter from being the method of choice, our result yields insights into the nature of the Kaiser window and its potential usefulness as a QPE-based alternative to QETU and other QSP-based filtering approaches.
In relation to this study, we note that the Chebyshev expansion coefficients $c_{2l}$ are sometimes calculated via the procedure proposed by Low and Chuang [49]. Their procedure approximates the discontinuous sign function by the error function, which is in turn approximated by using the known polynomial expansion of the exponential decay function. The corresponding expansion coefficients are obtained by evaluating the modified Bessel functions of the first kind [49]. We also estimate the number of degrees in this procedure using the analytic formula derived from it [50, 51], and find that it generally requires larger number of queries compared to the current convex-optimization-based results. This is understandable, as the convex-optimization-based approach does not assume any specific form for the polynomial function, therefore the expansion coefficients ($c_{l}$) are subject to fewer restrictions. However, getting the coefficients via numerical optimization becomes more difficult as one requires higher accuracy, therefore for very small $\delta$ or $\epsilon$ one has to rely on non-optimization-based approaches such as Low and Chuang’s procedure. In such cases, the QPE-based filter with the Kaiser window may be a competitive alternative.
@@ III.1 QPE-based spectral simulation
Finally, we consider an application of the present approach to the QPE-based calculation of the dynamical response function [34, 29, 30, 31, 32, 33, 35]. The objective is to calculate
@@
The final frequency resolution of this approach is $\Delta\omega^{\prime}=\frac{2\pi}{N^{\prime}T^{\prime}}$, with $N^{\prime}=2^{n^{\prime}}$. The parameters $n$ and $T$ used in the first filtering step can be smaller than those used in the second step, as in the coarse QPE approach proposed in Ref. [15]. In a naive application of QAA without filtering, the total number of queries to $U_{HT}$ is $O(p_{0}^{-1/2}\beta N^{\prime})$, where $\beta=T^{\prime}/T$ and $p_{0}$ is the probability of measuring $\tilde{S}_{B}(\omega_{y})$ in the energy range of interest. With filtering, the total number of queries is reduced to $O(p_{0}^{-1/2}N+\beta N^{\prime})$, with $N=O(\delta^{-1}\log\epsilon^{-1})$ for the case of Kaiser window-based filter. This two-step approach therefore allows for the efficient calculation of the energy spectrum of the specified small energy region when $\beta N^{\prime}\gg N$ and can be useful to get fine structures of the spectrum at high resolution.
@@ III.2 Density of states of MnO
To see the performance of the proposed QPE-based filtering for realistic spectra, we apply it to the calculation of the DOS [36] of the type-II antiferromagnetic MnO. This material is a prototype of transition metal monoxides, where strong electron correlation between the $3d$ orbitals of the transition metal atoms in a narrow energy region plays a crucial role in determining their electronic and magnetic properties. The insulating nature of the type-II antiferromagnetic MnO is explained by the combination of the magnetic ordering of Mn and the coupling between Mn $3d$ and O $2p$ orbitals [52, 53], and a number of band-structure calculations have been performed to understand the exact electronic structure of this system and also as a benchmark of a newly proposed algorithm  [54, 55, 56, 57, 58, 59, 60].
@@
where $\varepsilon_{\mathbf{k}\nu}$ are one-particle orbital energies, and $\mathbf{k}$ and $\nu$ are the Bloch wavevectors and orbital (or band) indices [36], respectively. Although this DOS itself can be calculated efficiently on a classical computer, we use it as an approximation of the spectral function for interacting many-electron systems, which is hard to calculate classically [30, 32, 31]. The DOS and spectral function are related to the experimental spectra obtained in photoemission spectroscopy.
In this work we prepare $\varepsilon_{\mathbf{k}\nu}=E_{\mu=(\mathbf{k}\nu)}$ classically using density functional theory [61] and set
$B=\sum_{\mathbf{k}\nu}\hat{c}_{\mathbf{k}\nu}^{\dagger}\hat{c}_{\mathbf{k}\nu},$ (57)
where $\hat{c}_{\mathbf{k}\nu}^{\dagger}$ and $\hat{c}_{\mathbf{k}\nu}$ are the electron creation and annihilation operators, respectively. The DOS is calculated via Eq. (55) using a noiseless simulator. The input data is prepared using QuantumEspresso [62, 63] with the pseudopotentials [36] from its pseudopotential database [64]. The rocksalt structure is assumed with lattice constant $a=4.43$. The generalized gradient approximation (GGA) [65] is used for the exchange and correlation functional, and the kinetic energy cutoff for wavefunctions used is $47$ Ry. The one-particle energies $\varepsilon_{\mathbf{k}\nu}$ are calculated with $32\times 32\times 32$ $\mathbf{k}$-points after obtaining the electron charge density self-consistently. Although the number of states in this calculation is prohibitively large for early fault-tolerant quantum computers, the resulting DOS yields a continuous spectrum that helps us to evaluate the efficacy of the proposed approach.
@@
The spectrum above the cutoff consists mainly of Mn $3d$ states. As can be seen in Fig. 11(b), due to the crystal field effect it splits into two main peaks [53, 52]: the lower-lying $t_{2g}$ peak and the higher-lying $e_{g}$ peak. The $e_{g}$ peak lies approximately from $0.42$ to $0.46$ hartree, and we set this energy region as the target energy range. We add a padding of 0.5 hartree above the $e_{g}$ states. Figures 11(b) and 11(c) show the DOS (obtained via Eq. (55)) after QPE-based filtering in open circles, calculated with the rectangular window (Fig. 11(b)) and with the Kaiser window using $\alpha=3$ (Fig. 11(c)). Note in both cases the sine window function is used in the second step. The parameters used are $n=n^{\prime}=6$ ($N=N^{\prime}=128$) and $T=2,T^{\prime}=20$, and the coarse QPE frequency grid points used in the filtering step are shown by the filled squares.
@@
The error is largest in the rectangular window result, and the error in the Kaiser window result decreases after ±$\pm\alpha$ frequency grid points used in filtering from the cutoff frequency $\omega_{y_{c}}$, in accordance with the observation in the previous section. Switching from the rectangular window to the Kaiser window reduces the error in the target energy region about $0.42$ to $0.46$ hartree, which corresponds to the $e_{g}$ orbital energies (see Figs. 11(b) and 11(c)), by about six orders of magnitude.
@@
In summary, this calculation demonstrates that the proposed two-step approach provides an efficient way of calculating high-resolution spectra of the specified small energy region. This approach is applicable to general discrete and continuous spectral simulations of quantum many-body systems, which are hard to calculate via classical means. The Kaiser window-based filter provides an accurate and controllable approach, with tunable parameters $N$ and $\alpha$ based on the specified target energy interval and the tolerance error $\epsilon$ using Eqs. (26) and (27).
@@
In this work, we have investigated the properties of the QPE-based filtering approach for three different window functions. We have shown that the results with the conventional rectangular window exhibit an oscillation known as the Gibbs oscillation, and this oscillation is suppressed when using the sine and the Kaiser windows. We have also shown that this difference is explained by the $\sigma$-factor (smoothing factor) appearing in the theory of the Gibbs oscillation, the tuning of which may contribute to the future improvement of window functions for QPE.
We have found that, under the same conditions the number of queries to $U_{HT}$ required in the Kaiser window-based filter is comparable to that in QETU after optimization of phase angles. This indicates that the form of the Kaiser window function is nearly optimal for filtering and can be used as a potential alternative to QETU, although one should keep in mind that there is an additional cost associated with window state preparation.
As an application of the current approach, we have considered the combination of the filtering technique to the QPE-based spectral calculation. We have also performed a numerical simulation of the DOS of type-II antiferromagnetic MnO as an approximation of the spectral function. Our results have shown that the QPE-based filtering efficiently extracts the energy range of interest, in this case the $e_{g}$ orbitals of Mn. This calculation has also shown that even the simplest rectangular window sufficiently reproduces the peak structures of this system. These results indicate that, combined with efficient ground-state preparation and time evolution techniques, the proposed algorithm can be used as a simple QPE-based scheme complementary to QSP-based approaches for future materials simulations on a fault-tolerant quantum computer.

### v3

@@ II.4 The Gibbs phenomenon and the $\sigma$-factor
As seen in the previous section, the QPE-based filter $\mathcal{R}_{\mu}$ generally exhibits oscillation known as the Gibbs phenomenon [Jerri1998]. To analyze this effect, let us first rewrite $\mathcal{R}_{\mu}$ as
@@
In the theory of the Gibbs phenomenon [Jerri1998], this $\sigma$-factor is called a “filter” (not to be confused with filtering of states discussed in earlier sections) and serves as a smoothing factor that improves the convergence of the truncated Fourier series. Several filters have been proposed in the classical Fourier spectral analysis. Some examples are the Lanczos filter and the raised cosine filter [Jerri1998, Vandeven1991].
@@ II.5 Comparison with QETU
[…]
@@
where $T_{2l}(w)$ and $c_{2l}$ are the $2l$-th Chebyshev polynomials and expansion coefficients, respectively. We numerically determine the smallest $d$ in which $F_{d}(E)=|f_{d}(\textrm{cos}\frac{ET}{4})|^{2}$ satisfies Eqs. (42-44) for all $E\in[0,\frac{2\pi}{T}]$.
We compare two approaches to determine $c_{2l}$: (1) the polynomial approximation of the erf function discussed by Low and Chuang [low2017hamiltoniansimulationuniformspectral] and (2) the convex optimization approach [PRXQuantum.3.040305].
The first approach uses the polynomial approximation of the erf function $\textrm{erf}(kx)(k>0)$ [low2017hamiltoniansimulationuniformspectral] that is also used in the previous QPE-QETU comparison in Ref. [PRXQuantum.5.040339]:
$p_{\textrm{erf},k,d+1}(x)$ $=$ $\frac{2ke^{-k^{2}/2}}{\sqrt{\pi}}\Bigl(I_{0}(k^{2}/2)x+$ (52)
$\sum_{j=1}^{d/2}I_{j}(k^{2}/2)(-1)^{j}$
$\bigl(\frac{T_{2j+1}(x)}{2j+1}-\frac{T_{2j-1}(x)}{2j-1}\bigr)\Bigr).$
The function $f_{d}(w)$ is constructed following the approximation of the rectangular function [low2017hamiltoniansimulationuniformspectral]
$f_{d}(w)$ $=$ $c\Bigl[1-\frac{1}{2}\Bigl(p_{\textrm{erf},2k,d+1}((w+w_{m})/2)$ (53)
$+p_{\textrm{erf},2k,d+1}((-w+w_{m})/2)\Bigr)\Bigr]$
$w_{m}$ $=$ $\frac{w_{+}+w_{-}}{2}$ (54)
$k$ $=$ $\frac{\sqrt{2}}{w_{+}-w_{-}}\log^{1/2}\bigl(\frac{2}{\pi\tilde{\epsilon}^{2}}\bigr)$ (55)
$\tilde{\epsilon}$ $=$ $1-\sqrt{1-\epsilon}.$ (56)
This $f_{d}(w)$ $\tilde{\epsilon}-$approximates $f(w)$ with $d=O(\frac{1}{w_{+}-w_{-}}\log\tilde{\epsilon}^{-1})=O(\frac{1}{\delta}\log\epsilon^{-1})$ [low2017hamiltoniansimulationuniformspectral].
In the second approach, the expansion coefficients $c_{2l}$ are calculated with the convex-optimization-based approach proposed in Ref. [PRXQuantum.3.040305] and using the qsppack library [PhysRevA.103.042419, Wang2022energylandscapeof, Dong2024infinitequantum]. We use $M$ grid points $w_{m=0,1,\dots,M-1}$ in $[0,1]$ and solve the following optimization problem
@@
In both approaches, we check whether the function $F_{d}(E)$ satisfies Eqs. (42-44) for all $E\in[0,\frac{2\pi}{T}]$ by combining a grid search with a fine grid of $100,000$ points and an optimization routine using the scipy.optimization module [2020SciPy-NMeth]. As an example of the calculated results, $f_{d}(w)$ and $F_{d}(E)$ obtained for $\epsilon=0.01$ are shown in Figs. 8(a) and 8(b).
@@
Figure 8: Polynomial approximations of (a) $f(w)$ and (b) $F(E)=|f(\cos\frac{ET}{4})|^{2}$ for $\epsilon=0.01$ calculated with the prescription of Low and Chuang [low2017hamiltoniansimulationuniformspectral] with degree 1226 and via the convex optimization [PRXQuantum.3.040305] with degree 144. The parameters used are $E_{\textrm{targ}}=1.4430,2\delta=0.2556$, and $T=1$. The two vertical dotted lines indicate the positions of $w_{\pm}$ in (a) and $E_{\textrm{targ}}$ and $E_{\textrm{targ}}+2\delta$ in (b). The inset in (b) shows a magnified view.
@@
Figure 9: Minimum number of queries to $U_{HT}$ in QETU for three sets of ($E_{\textrm{targ}},\delta$). The corresponding values in the QPE-based filters are also shown for comparison. The straight lines show the results of the linear regression. (a) $E_{\textrm{targ}}=1.0589$ and $2\delta=1.0232$. (b) $E_{\textrm{targ}}=1.4430$ and $2\delta=0.2556$. (c) $E_{\textrm{targ}}=1.5069$ and $2\delta=0.1279$. In (c) the results for $\epsilon=10^{-4}$ and $\epsilon=10^{-5}$ with $M=800$ are not calculated as there are less sampled points than the required degree $d$.
@@
Figure 9 shows the calculated number of queries to $U_{HT}$, which is half of the minimum degree $d$, for three different sets of $(E_{\textrm{targ}},\delta)$. Here $E_{\textrm{targ}}$ and $\delta$ are obtained numerically from the results of the Kaiser window-based filter (Fig. 5 (b)) with $\alpha=3$ and $n=6,8,9$, using $y_{c}=2^{n-2}-1$ and $\epsilon=10^{-7}$ (see also Fig. 4). The number of sampled points $M$ in the second approach is a hyperparameter, and we compare three cases, $M=800,3200$ and $12800$. In the figures, $\epsilon$ is varied from $10^{-1}$ to $10^{-2}$ ($10^{-2}$ to $10^{-5}$) for the first (second) approach. The errors in the QPE-based filters with the three window functions are also shown for comparison.
It can be seen that the number of queries in the first approach is noticeably larger than that in the second approach; for $\epsilon=10^{-2}$, queries in the former approach are $\approx 9$ times larger than those in the latter. As a result, the QETU-based filter in the first approach requires more queries than the QPE-based filters. This aligns with the conclusions drawn in Ref. [PRXQuantum.5.040339] based on the rectangular window, which state that the QPE-based filtering is more efficient than QETU.
@@
The number of required queries in the second approach, on the other hand, is smaller than those in the QPE-based filters with the rectangular and sine windows. It decreases as the number of sampled points $M$ increases, which can be understood by noting that with a small $M$ the optimized coefficients tend to overfit to the sampled data points. Interestingly, for the three cases studied, the QETU results in the second approach extrapolated to $\epsilon=10^{-7}$ are very close to the results of the Kaiser window-based filter (crosses). Although the required queries depend on the form of the QETU filter function used, this result can partly be understood by noting that the Kaiser window is an approximation of the prolate spheroidal window, which is designed to maximize the probability of the mainlobe of the probability distribution $P_{\mu}(y)$ [patel2024optimalcoherentquantumphase]. This construction is approximately equivalent to minimizing the error of the QPE-based filter $\Delta\mathcal{R}_{\mu}$ defined in Eq. (14), potentially explaining the similarities between the QETU-based filter obtained from the optimization (Eq. (57)) and the Kaiser window-based filter. We note that the present calculation compares two slightly different filter forms, but our comparison provides insight into the Kaiser window-based filter as an alternative to QETU and other QSP-based filtering methods.
We next discuss two situations in which the QPE-based filter could be advantageous. The first case is when we have prior knowledge of the input spectrum. For example, when we know the lower bound of the input spectrum $E_{\textrm{min}}$ satisfying $\delta<E_{\textrm{min}}$, we can reduce the error $\Delta\mathcal{R}_{\mu}$ for the states with $E_{\mu}\in[\frac{2\pi}{T}-\frac{\delta}{2},\frac{2\pi}{T}]$ (see Fig. 5(b)) by shifting the Hamiltonian and the cutoff frequency as $H\to H-\frac{\delta}{2}$ and $\omega_{y_{c}}\to\omega_{y_{c}}-\frac{\delta}{2}$, respectively. For this shifted system, $\mathcal{R}_{\mu}$ for the Kaiser window-based filter satisfies
$1-\epsilon$ $\leq$ $\mathcal{R}_{\mu}\quad\quad E_{\mu}^{\prime}\in\left[0,E_{\textrm{targ}}\right],$ (58)
$\mathcal{R}_{\mu}$ $\leq$ $\epsilon\quad\quad E_{\mu}^{\prime}\in\left[E_{\textrm{targ}}+\delta,\frac{2\pi}{T}\right],$ (59)
where $E_{\textrm{targ}}$ is defined in Eq. (27) and $E_{\mu}^{\prime}$ are shifted eigenenergies. These two conditions are identical to the QETU ones given in Eqs. (43) and (44) except that $2\delta$ in Eq. (44) is replaced by $\delta$. Since the number of queries in QETU scales as $O(\delta^{-1}\log\epsilon^{-1})$, this replacement approximately doubles the number of necessary queries in QETU. Therefore, in this case the Kaiser window-based filter is expected to be more efficient than the optimization-based QETU filter. A similar argument can be applied when the upper bound of the input spectrum $E_{\textrm{max}}$ is known. Furthermore, no shifting is required if both $E_{\textrm{min}}$ and $E_{\textrm{max}}$ are known and satisfy $E_{\textrm{min}}>\frac{\delta}{2}$ and $E_{\textrm{max}}<\frac{2\pi}{T}-\frac{\delta}{2}$, respectively.
The second case is when a numerical optimization procedure cannot be performed as, for example, the required accuracy is very high; in our calculation we have difficulty obtaining reliable minimum degrees for $\epsilon<10^{-5}$. For very small $\delta$ and $\epsilon$ one may have to rely on simpler non-optimization-based approaches such as Low and Chuang’s procedure. In such cases, the QPE-based filter with the Kaiser window can be a competitive alternative to QETU.
Note that the QPE-based approach requires $n=O\Bigl(\log\bigl[\delta^{-1}\log\epsilon^{-1}\bigr]\Bigr)$ ancilla qubits, and there are additional costs from the preparation of the Kaiser window as well as QFT. The gate complexity of QFT is $O(n^{2})$, and as discussed in Appendix B, the gate complexity of the QSVT-based approach in Ref. [mcardle2022quantumstatepreparationcoherent] is $O\Bigl(n\sqrt[4]{\pi\alpha+1}(\pi\alpha+\log\epsilon^{-1})\Bigr)=$ $O\Bigl(\log\Bigl[\delta^{-1}\log\epsilon^{-1}\Bigr](\log\epsilon^{-1})^{5/4}\Bigr)$ with three extra ancilla qubits. These costs are exponentially smaller than that of $N=2^{n}$ queries to $U_{HT}$ and $U_{HT}^{\dagger}$, therefore they do not become the bottleneck of the calculation for large systems. We also note that QETU requires $O(N)$ single-qubit rotations (see Fig. 7), and this cost is absent in the QPE-based filtering.
Finally, we briefly discuss general errors in practical calculations for both QPE and QETU approaches, which are ignored in the simulation. The most important error comes from time propagation. Using Trotterization [PhysRevX.11.011020], the time propagation operator $U_{HT}$ is approximated as $U_{HT}=\Bigl(\exp[iHT/r]\Bigr)^{r}\approx S_{p}(T/r)^{r}$, where $r$ is the Trotter step, and $S_{p}(t)$ is the $p$-th order Trotter product. The Trotter error can be evaluated in the operator norm
$U_{HT}-S_{p}(T/r)^{r}||=O(C_{\textrm{Trotter}}r^{-p}),$ (60)
where $C_{\textrm{Trotter}}$ is a system- and $p$-dependent constant factor constructed from nested commutators [PhysRevX.11.011020]. A number of works have investigated tighter, more practical bounds [10.5555/2871401.2871402, PhysRevA.91.022311, Su2021nearlytight, PhysRevA.105.012403, Sahinoglu2021, Chen2024, PhysRevLett.129.270502, blunt2025montecarloapproachbound]. With $N$ queries to $U_{HT}$ and $U_{HT}^{\dagger}$, the required steps for precision $\epsilon_{\textrm{Trotter}}$ is $r=O(N^{\frac{1}{p}}C_{\textrm{Trotter}}^{\frac{1}{p}}\epsilon_{\textrm{Trotter}}^{-\frac{1}{p}})$ [PRXQuantum.3.040305]. To avoid the Trotter error, qubitization and quantum walker based approaches have been proposed [Low2019hamiltonian, Berry2018], and they can also be combined with the current filtering.
In the early fault-tolerant era the imperfectness of quantum hardware also affects the results. Various quantum error mitigation techniques have to been proposed to remove errors for expectation value estimation and for sampling [RevModPhys.95.045005, liu2025quantumerrormitigationsampling], and these techniques can be applied to obtain correct outputs.
@@ III.1 QPE-based spectral simulation
Finally, we consider an application of the present approach to the QPE-based calculation of the dynamical response function [PhysRevA.92.062318, PhysRevC.100.034610, PhysRevA.101.012330, PhysRevA.110.022618, PhysRevA.110.L060401, fomichev2024simulatingxrayabsorptionspectroscopy, nishi2025demonstrationlogicalquantumphase]. The objective is to calculate
@@
The final frequency resolution of this approach is $\Delta\omega^{\prime}=\frac{2\pi}{N^{\prime}T^{\prime}}$, with $N^{\prime}=2^{n^{\prime}}$. The parameters $n$ and $T$ used in the first filtering step can be smaller than those used in the second step, as in the coarse QPE approach proposed in Ref. [PRXQuantum.5.040339]. In a naive application of QAA without filtering, the total number of queries to $U_{HT}$ is $O(p_{0}^{-1/2}\beta N^{\prime})$, where $\beta=T^{\prime}/T$ and $p_{0}$ is the probability of measuring $\tilde{S}_{B}(\omega_{y})$ in the energy range of interest. With filtering, the total number of queries is reduced to $O(p_{0}^{-1/2}N+\beta N^{\prime})$, with $N=O(\delta^{-1}\log\epsilon^{-1})$ for the case of Kaiser window-based filter. This two-step approach therefore allows for the efficient calculation of the energy spectrum of the specified small energy region when $\beta N^{\prime}\gg N$ and can be useful to get fine structures of the spectrum at high resolution.
@@ III.2 Density of states of MnO
To see the performance of the proposed QPE-based filtering for realistic spectra, we apply it to the calculation of the DOS [martin2020electronic] of the type-II antiferromagnetic MnO. This material is a prototype of transition metal monoxides, where strong electron correlation between the $3d$ orbitals of the transition metal atoms in a narrow energy region plays a crucial role in determining their electronic and magnetic properties. The insulating nature of the type-II antiferromagnetic MnO is explained by the combination of the magnetic ordering of Mn and the coupling between Mn $3d$ and O $2p$ orbitals [PhysRevB.30.4734, PhysRevLett.52.1830], and a number of band-structure calculations have been performed to understand the exact electronic structure of this system and also as a benchmark of a newly proposed algorithm  [PhysRevB.51.1477, PhysRevB.76.165106, PhysRevB.78.155112, PhysRevLett.103.036404, PhysRevB.79.235114, PhysRevX.5.011006, PhysRevB.101.165138].
@@
where $\varepsilon_{\mathbf{k}\nu}$ are one-particle orbital energies, and $\mathbf{k}$ and $\nu$ are the Bloch wavevectors and orbital (or band) indices [martin2020electronic], respectively. Although this DOS itself can be calculated efficiently on a classical computer, we use it as an approximation of the spectral function for interacting many-electron systems, which is hard to calculate classically [PhysRevA.101.012330, PhysRevA.110.L060401, PhysRevA.110.022618]. The DOS and spectral function are related to the experimental spectra obtained in photoemission spectroscopy.
In this work we prepare $\varepsilon_{\mathbf{k}\nu}=E_{\mu=(\mathbf{k}\nu)}$ classically using density functional theory [dreizler2012density] and set
$B=\sum_{\mathbf{k}\nu}\hat{c}_{\mathbf{k}\nu},$ (65)
where $\hat{c}_{\mathbf{k}\nu}$ are the electron annihilation operators. The DOS is calculated via Eq. (63) using a noiseless simulator. The input data is prepared using QuantumEspresso [QE-2017, QE-2009] with the pseudopotentials [martin2020electronic] from its pseudopotential database [QE-PPDB]. The rocksalt structure is assumed with lattice constant $a=4.43$ Å. The generalized gradient approximation (GGA) [PhysRevLett.77.3865] is used for the exchange and correlation functional, and the kinetic energy cutoff for wavefunctions used is $47$ Ry. The one-particle energies $\varepsilon_{\mathbf{k}\nu}$ are calculated with $32\times 32\times 32$ $\mathbf{k}$-points after obtaining the electron charge density self-consistently. Although the number of states in this calculation is prohibitively large for early fault-tolerant quantum computers, the resulting DOS yields a continuous spectrum that helps us to evaluate the efficacy of the proposed approach.
@@
The spectrum above the cutoff consists mainly of Mn $3d$ states. As can be seen in Fig. 11(b), due to the crystal field effect it splits into two main peaks [PhysRevLett.52.1830, PhysRevB.30.4734]: the lower-lying $t_{2g}$ peak and the higher-lying $e_{g}$ peak. The $e_{g}$ peak lies approximately from $0.42$ to $0.46$ hartree, and we set this energy region as the target energy range. We add a padding of 0.5 hartree above the $e_{g}$ states. Figures 11(b) and 11(c) show the DOS (obtained via Eq. (63)) after QPE-based filtering in open circles, calculated with the rectangular window (Fig. 11(b)) and with the Kaiser window using $\alpha=3$ (Fig. 11(c)). Note in both cases the sine window function is used in the second step. The parameters used are $n=n^{\prime}=6$ ($N=N^{\prime}=128$) and $T=2,T^{\prime}=20$, and the coarse QPE frequency grid points used in the filtering step are shown by the filled squares.
@@
The error is largest in the rectangular window result, and the error in the Kaiser window result decreases after $\pm\alpha$ frequency grid points used in filtering from the cutoff frequency $\omega_{y_{c}}$, in accordance with the observation in the previous section. Switching from the rectangular window to the Kaiser window reduces the error in the target energy region about $0.42$ to $0.46$ hartree, which corresponds to the $e_{g}$ orbital energies (see Figs. 11(b) and 11(c)), by about six orders of magnitude.
@@
In summary, this calculation demonstrates that the proposed two-step approach provides an efficient way of calculating high-resolution spectra of the specified small energy region. This approach is applicable to general discrete and continuous spectral simulations of quantum many-body systems, which are hard to calculate via classical means. Our results show that the simplest rectangular window-based filter already reproduces the main peak structures of the system; as the rectangular and sine windows are easy to construct, they can prove useful when the required accuracy is not very high. The Kaiser window-based filter provides an accurate and controllable approach, with tunable parameters $N$ and $\alpha$ based on the specified target energy interval and the tolerance error $\epsilon$ using Eqs. (26) and (27).
@@
In this work we have investigated the properties of the QPE-based filtering approach and have shown that the choice of window functions significantly affects the quality of the results. The results with the conventional rectangular window exhibit an oscillation known as the Gibbs oscillation, which originates from large spectral leakage for this window. This oscillation is suppressed when using the sine and the Kaiser windows, and this difference is explained by the $\sigma$-factor (smoothing factor) appearing in the analysis of the Gibbs oscillation. Further analysis and tuning of this factor may provide a way of improving window functions for QPE.
We have also compared the performance of the QPE-based filter and QETU in terms of necessary queries to the time propagation operator. We have found that, under the same conditions, the number of queries required in the Kaiser window-based filter is comparable to that in QETU after optimization of phase angles. This indicates that the form of the Kaiser window function is nearly optimal for filtering and can be used as a potential alternative to QETU. We have also estimated the cost of window preparation and have shown that it is exponentially smaller than that of the time propagation operations. We have also discussed that the Kaiser window-based filter can be more cost-effective than QETU when we have prior knowledge of the energy spectrum or filter construction in QETU is computationally hard.
As an application of the current approach, we have considered the combination of the filtering technique with the QPE-based spectral calculation. To assess the performance of the QPE-based filters for realistic spectra in materials science, we have performed a numerical simulation of the DOS of type-II antiferromagnetic MnO as an approximation of the spectral function. Our results have shown that the QPE-based filtering efficiently extracts the states in the energy range of interest, in this case the $e_{g}$ orbitals of Mn. Even the simplest rectangular window sufficiently reproduces the peak structures of this system, and the results can further be refined in a controllable manner by using the Kaiser window. These results indicate that, combined with efficient ground-state preparation and time evolution techniques, the proposed algorithm can be used as a simple QPE-based scheme complementary to QSP-based approaches for future materials simulations on a fault-tolerant quantum computer.